Hi,
Could u pls help me with this:
Find a number that is between : 3/7 & 5/11
Thanks a lot,
H.H

Answer by Alan3354(69443)

Find a number that is between : 3/7 & 5/11
------------
0.43
-----
If you prefer a fraction, find a common denominator.
7*11 = 77
Use a multiple of 77, say 7700
--
3/7 = 3300/7700
5/11 = 3500/7700
-------
Any numerator between 3300 & 3500 / 7700 is between
3301, 3302, 3302.56689, etc